IGO Night Logs Cycle 2010A

April 2010 (Total available observing time: 300 hours)

Hours Observed: 167 hours
Weather Loss: 46 hours
Technical Loss: 81 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Apr 30

0


10


Cassegrain Rotator Problem.

Apr 29

0


10


Cassegrain Rotator Problem.

Apr 28

0


10


Cassegrain Rotator Problem.

Apr 27

0


10


Cassegrain Rotator Problem.

Apr 26

0


10


Cassegrain Rotator Problem.

Apr 25

8

0

2

0

Clear sky. Cass Rotator problem. Observations carried out.

Apr 24

4

5

1

0

Cloudy sky. OPER ON had to be given. A few observations were carried out.

Apr 23

9

0

1

0

Passing clouds. Oil chiller fault. Observations carried out.

Apr 22

0

0

10

0

Polarimetry wheel got stuck in IFOSC. No observations.

Apr 21

7

0

1

2

Clear Sky. Observations carried out.

Apr 20

10

0

0

0

Clear Sky. Observations carried out.

Apr 19

9

0

1

0

Clear Sky. Observations carried out. AZM L2 limit occurred. Problem solved.

Apr 18

10

0

0

0

Clear Sky. Observations carried out. CCD warming up in the morning.

Apr 17

4

5

1

0

Cloudy sky. Observations taken through passing clouds. Primary Mirror support problem.

Apr 16

5

5

0

0

Cloudy sky. Observations taken through passing clouds.

Apr 15

0

0

10

0

Primary Mirror support problem. No observations.

Apr 14

9

0

0

1

Clear Sky. Observations carried out.

Apr 13

9

0

1

0

Clear Sky. Observations carried out. CCD warming up in the morning.

Apr 12

8

1

1

0

Partially cloudy sky. Observations carried out. CCD warming up in the morning.

Apr 11

9

0

1

0

Clear Sky. Observations carried out. CCD warming up in the morning.

Apr 10

7

3

0

0

Cloudy sky. Observations carried out.

Apr 09

3

7

0

0

Cloudy sky. A few observations taken through passing clouds.

Apr 08

0

10

0

0

Cloudy sky. No observations.

Apr 07

0

10

0

0

Cloudy sky. No observations.

Apr 06

10

0

0

0

Absorbed sky. Observations carried out.

Apr 05

7

0

0

3

Clear Sky. Observations carried out.

Apr 04

10

0

0

0

Clear Sky. Observations carried out.

Apr 03

9

0

1

0

Clear Sky. Observations carried out. PMS and AGS problem, solved.

Apr 02

10

0

0

0

Clear Sky. Observations carried out.

Apr 01

10

0

0

0

Clear Sky. Observations carried out.



March 2010 (Total available observing time: 341 hours)

Hours Observed: 235 hours
Weather Loss: 83 hours
Technical Loss: 11 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Mar 31

6

0

0

5

Clear sky. Observations carried out.

Mar 30

11

0

0

0

Clear sky. Observations carried out.

Mar 29

10

0

0

1

Absorbed sky. Pointing model applied. Observations carried out.

Mar 28

6

4

1

0

Passing clouds. Data collected for pointing model. Azimuth L2 limit occurred during the night.

Mar 27

10

0

1

0

Clear sky. Pointing Model generated. Azimuth L2 limit occurred during the night.

Mar 26

10

0

1

0

Clear Sky. Observations carried out. CCD became hot towards morning.

Mar 25

11

0

0

0

Clear Sky. Observations carried out.

Mar 24

11

0

0

0

Clear Sky. Observations carried out.

Mar 23

11

0

0

0

Clear Sky. Observations carried out.

Mar 22

11

0

0

0

Partially clear sky. Observations carried out.

Mar 21

7

4

0

0

Moving clouds. Observations carried out.

Mar 20

7

3

1

0

Partially clear sky. Observations carried out. Altitude axis failed.

Mar 19

7

4

0

0

Partially clear sky. Observations carried out.

Mar 18

3

7

1

0

Partially Clear sky. Observations carried out. Oil Chiller fault.

Mar 17

10

0

1

0

Partially Clear sky. Observations carried out. CCD warmed up in the morning.

Mar 16

0

11

0

0

Cloudy sky. No observations.

Mar 15

11

0

0

0

Partially Clear sky. Observations carried out.

Mar 14

8

3

0

0

Passing Clouds. Observations carried out.

Mar 13

0

11

0

0

Cloudy sky. No observations.

Mar 12

0

11

0

0

Clouds and rain. No observations.

Mar 11

0

11

0

0

Overcast. No observations.

Mar 10

1

10

0

0

Clouds and rain. Only a few frames taken.

Mar 09

6

3

2

0

Cloudy sky. Oil Chiller fault. Observations carried out.

Mar 08

11

0

0

0

Clear sky. Observations carried out.

Mar 07

10

0

1

0

Clear sky. Autoguider system failed. Had to be rebooted. Observations carried out.

Mar 06

11

0

0

0

Clear sky. Observations carried out.

Mar 05

10

0

1

0

Clear sky. AGF problem. Observations carried out.

Mar 04

9

1

1

0

Cloudy sky. Telescope focus problem. Observations carried out.

Mar 03

11

0

0

0

Data collected for pointing model.

Mar 02

6

0

0

5

Cloudy sky. Data collected for pointing model. GRB observations carried out.

Mar 01

10

0

0

1

Thin Clouds. Observations carried out.



February 2010 (Total available observing time: 308 hours)

Hours Observed: 221 hours
Weather Loss: 60 hours
Technical Loss: 07 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Feb 28

6

5

0

0

Cloudy sky. Observations carried out.

Feb 27

10

0

1

0

Clear sky. Telescope took auto shut down. Observations carried out.

Feb 26

10

0

1

0

Clear sky Autoguider Deployment Problem. Observations carried out

Feb 25

10

0

1

0

Thin clouds. Autoguider Focus Problem. Observations carried out

Feb 24

3

0

0

8

Thin clouds. Alignment of auto guider and IFOSC done.

Feb 23

8

0

3

0

Clear Sky. Autoguider Focus Problem. Observations carried out.

Feb 22

9

2

0

0

Moving clouds. Observations carried out.

Feb 21

0

11

0

0

Cloudy sky. No observations.

Feb 20

11

0

0

0

Thin clouds. Observations carried out.

Feb 19

11

0

0

0

Thin clouds. Observations carried out.

Feb 18

11

0

0

0

Thin clouds. Observations carried out.

Feb 17

10

0

0

1

Passing Clouds. Observations carried out.

Feb 16

4

5

0

2

Cloudy sky. Observations carried out.

Feb 15

0

11

0

0

Cloudy sky. No observations.

Feb 14

7

4

0

0

Cloudy sky. Observations carried out.

Feb 13

11

0

0

0

Clear sky. Observations carried out.

Feb 12

11

0

0

0

Clear sky. Observations carried out.

Feb 11

11

0

0

0

Passing clouds. Observations carried out.

Feb 10

0

11

0

0

Cloudy Sky. No observations.

Feb 09

8

3

0

0

Cloudy sky. Observations carried out

Feb 08

8

0

0

3

Clear sky. Observations carried out.

Feb 07

8

0

0

3

Clear sky. Observations carried out.

Feb 06

7

0

1

3

Clear sky. AGF problem occurred. Observations carried out.

Feb 05

11

0

0

0

Clear sky. Observations carried out.

Feb 04

10

1

0

0

Clear sky. Some problems with CCU software working.

Feb 03

11

0

0

0

Moving clouds

Feb 02

6

5

0

0

Clody sky.

Feb 01

9

2

0

0

Partial clouds.